OpenStudy (ttop0816):

is my function correct?? The sine function that has been reflected over the x-axis, has an amplitude of 2, and has a frequency of 6 -2sin6x <

OpenStudy (anonymous):

if "frequency" is the same as period, then no

OpenStudy (anonymous):

the \(-2\) is correct

OpenStudy (primeralph):

2pi*f = the coefficient of x. So, it's -2sin(12pi*x)

OpenStudy (ranga):

Reflected over the x axis means y should be replaced with -y. Or simply insert a negative sign before the sin(x). In f(x) = Asin(Bx), A is the amplitude and frequency = B/2pi

OpenStudy (anonymous):

\(\frac{2\pi}{b}=6\) gives \(b=\frac{\pi}{3}\)

OpenStudy (ranga):

-2sin(12pi*x)

OpenStudy (ranga):

Period = 2(pi)/B and frequency = 1 /period or B/2pi